OpenStudy (anonymous):

Which of the following is not equivalent to the formula for angular velocity, v = wr? vw = r v over w= r v over w r= 1 v over r= w

OpenStudy (paxpolaris):

\[\Large v=w \cdot r\]if we divide both sides by w, we get:\[\Large \frac vw=r\] instead, if we had divided both sides by (w r), we'd get:\[\Large {v \over w \cdot r}=1\] instead, if we had divided both sides by (w r), we'd get:\[\Large {v \over r}=w\]

OpenStudy (paxpolaris):

the only thing we can't get from v=wr is: vw=r

OpenStudy (kainui):

Wait a second. Can we divide by a vector?